OpenStudy (anonymous):

Simplify 6 square root of 44 end root minus square root of 45 end root minus square root 99 end root plus square root of 20

OpenStudy (anonymous):

Does it look like this? \[6\sqrt(44)-\sqrt(45)-\sqrt(99)+\sqrt(20)\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[6\sqrt(4*11)-\sqrt(9*5)-\sqrt(9*11)+\sqrt(4*5)\] \[6*2\sqrt(11)-3\sqrt(5)-3\sqrt(11)+2\sqrt(5)\] \[9\sqrt(11)-\sqrt(5)\]

OpenStudy (anonymous):

Thanks

OpenStudy (zzr0ck3r):

@DoYourHomework use {} with sqrt and it will look better \[\sqrt{this}\space vs\space\sqrt(this)\]

OpenStudy (anonymous):

Oh, sweet. Didn't know that. Thx.

OpenStudy (zzr0ck3r):

fur sure:) I just figured it out also